OpenStudy (anonymous):

A boy leaves on a bicycle trip at the rate of 8 mph. One hour later the father, realizing that his son forgot his camping gear, sets out by car. How fast must the father drive to overtake the boy in 15 minutes?

OpenStudy (anonymous):

Boy position equation= 8t+8 Father position equation= at (8t+8)=at

OpenStudy (anonymous):

T denote time in hour 15/160= 1/4 hours 8(1/4)+8=a(1/4) 2+8=1/4 a 10 = 1/4 a a= 40 mile per hours

OpenStudy (anonymous):

okay thanks